OpenStudy (anonymous):

Write 32x5y5 with only one exponent. Use parentheses.

OpenStudy (anonymous):

How would I do this?

OpenStudy (anonymous):

\[32x ^{5}y ^{5}\]

OpenStudy (ajprincess):

\(32=2^5\) so \(32x^5y^5=2^5x^5y^5\)

OpenStudy (ajprincess):

U can write the answer in this form \(a^b*c^b*d^b=(acd)^b\)

OpenStudy (ajprincess):

Does that help?

OpenStudy (anonymous):

Yes, thank you. So would that be \[(2xy)^{5}\]

OpenStudy (ajprincess):

ya:)

OpenStudy (anonymous):

Thank you very much

OpenStudy (ajprincess):

welcome:)